# Effects of short-range correlations at high densities on neutron stars with and without DM content: role of the repulsive self-interaction

### v2

@@ I Introduction
In conventional relativistic mean-field (RMF) models, the nucleon momentum distribution is typically described by a step function, corresponding to a free Fermi gas of independent particles. This approximation neglects the effects of nucleon-nucleon correlations at short distances, which have been experimentally observed and theoretically studied in recent years. Short-range correlations (SRC) modify this picture by introducing a high-momentum tail (HMT) in the distribution, reflecting the presence of correlated nucleon pairs with large relative momentum [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] . These correlations are supported by electron-induced nucleon knockout experiments, and can be implemented in RMF models through phenomenological parametrizations that replace the step function with a distribution proportional to $1/k^{4}$ above the Fermi surface. This modification has been shown to impact the thermodynamic properties of nuclear matter and the structure of neutron stars. Nonlinear RMF extensions, such as the inclusion of a quartic $\omega^{4}$ term, have been shown to regulate the stiffness of the EoS at high densities [20]. As demonstrated in Ref. [20], this term arises naturally in effective field theory and significantly affects neutron star mass predictions, even when models share identical saturation properties.
@@ II Relativistic short-range correlated model
$=\frac{\partial(\mathcal{E}_{\mathrm{kin},p}+\mathcal{E}_{\mathrm{kin},n})}{\partial n_{q}}\Bigg|_{\rho_{\bar{q}},W,R,\Phi}+\frac{\partial\mathcal{E}_{\mathrm{pot}}}{\partial n_{q}}\Bigg|_{\rho_{\bar{q}},W,R,\Phi},$ (22)
@@ III.3 Impact for neutron stars
This behavior is particularly relevant in light of recent astrophysical observations reporting compact stars with masses around or above two solar masses, which impose stringent constraints on viable nuclear models. To illustrate this case, Fig. 6b shows parameterizations where the fourth-order term in $W$ is present in the vector channel ($C\neq 0$). A clear increase in the stellar mass is observed across the entire range of central energy densities analyzed. Notably, this increase is more pronounced than the decrease seen in the $C=0$ scenario. In summary, the influence of SRC on NS properties is highly sensitive to the structure of the vector self-interaction in the model. While SRC tend to soften the EoS when $C=0$, they can induce significant stiffening for $C\neq 0$, thereby enabling consistency with observations of massive NSs.
It should be noted that Fig. 6 includes dynamically unstable stars ($\partial M/\partial\epsilon_{c}<0$) to emphasize that the asymptotic high-density behavior persists across the density range relevant for NSs. Omitting these segments would obscure this continuity and potentially mislead interpretation, so we retain them to clearly illustrate the impact of SRC throughout the entire domain.
@@ IV Consequences in neutron stars with DM content
Figure 10 illustrates how the mass of the last stable star, as determined from Figs. 8 and 9, varies as a function of the DM mass fraction $F_{\text{dm}}$.
@@ Acknowledgements.
This work has been done as a part of the Project INCT-Física Nuclear e Aplicações, under No. 408419/2024-5. It is also supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under Grants No. 307255/2023-9 (O.L.), No. 308528/2021-2 (M.D.), No. 01565/2023-8 (Universal - O.L, M.D.), No. 409736/2025-2 (Universal - O.L, M.D.), and Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) under Thematic Project No. 2024/17816-8 (O.L., M.D.), and Project No 2025/06216-2 (C. B.).

### v3

@@ I Introduction
In conventional relativistic mean-field (RMF) models, the nucleon momentum distribution is typically described by a step function, corresponding to a free Fermi gas of independent particles. This approximation neglects the effects of nucleon-nucleon correlations at short distances, which have been experimentally observed and theoretically studied in recent years. Short-range correlations (SRC) modify this picture by introducing a high-momentum tail (HMT) in the distribution, reflecting the presence of correlated nucleon pairs with large relative momentum [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, revsrc1, revsrc2] . These correlations are supported by electron-induced nucleon knockout experiments, and can be implemented in RMF models through phenomenological parametrizations that replace the step function with a distribution proportional to $1/k^{4}$ above the Fermi surface. This modification has been shown to impact the thermodynamic properties of nuclear matter and the structure of neutron stars. Nonlinear RMF extensions, such as the inclusion of a quartic $\omega^{4}$ term, have been shown to regulate the stiffness of the EoS at high densities [20]. As demonstrated in Ref. [20], this term arises naturally in effective field theory and significantly affects neutron star mass predictions, even when models share identical saturation properties.
@@ II Relativistic short-range correlated model
$=\frac{\partial(\mathcal{E}_{\mathrm{kin},p}+\mathcal{E}_{\mathrm{kin},n})}{\partial\rho_{q}}\Bigg|_{\rho_{\bar{q}},W,R,\Phi}+\frac{\partial\mathcal{E}_{\mathrm{pot}}}{\partial\rho_{q}}\Bigg|_{\rho_{\bar{q}},W,R,\Phi},$ (22)
@@ III.3 Impact for neutron stars
This behavior is particularly relevant in light of recent astrophysical observations reporting compact stars with masses around or above two solar masses, which impose stringent constraints on viable nuclear models. To illustrate this case, Fig. 6b shows parameterizations where the fourth-order term in $W$ is present in the vector channel ($C\neq 0$). A clear increase in the stellar mass is observed across the entire range of central energy densities analyzed. Notably, this increase is more pronounced than the decrease seen in the $C=0$ scenario. In summary, the influence of SRC on NS properties is highly sensitive to the structure of the vector self-interaction in the model. While SRC tend to soften the EoS when $C=0$, they can induce significant stiffening for $C\neq 0$, thereby enabling consistency with observations of massive NSs. For the sake of completeness, in Table 2 we show the maximum mass values obtained for each parametrization shown in Fig. 6, in order to quantify the effects induced by SRC in each type of model.
Table 2: Maximum stellar mass, in units of $M_{\odot}$, for the mass-radius diagrams presented in Fig. 6.
models $M_{\mathrm{max}}$ $M^{\mbox{\tiny SRC}}_{\mathrm{max}}$
$C=0$
NL3*/NL3*-SRC $2.76$ $2.73$
NLM/NLM-SRC $2.27$ $2.21$
GM3/GM3-SRC $2.02$ $1.95$
$C\neq 0$
PK1/PK1-SRC $2.32$ $2.41$
TMA/TMA-SRC $1.99$ $2.14$
MB/MB-SRC $1.55$ $1.76$
It should be noted that Fig. 6 includes dynamically unstable stars ($\partial M/\partial\mathcal{E}_{c}<0$) to emphasize that the asymptotic high-density behavior persists across the density range relevant for NSs. Omitting these segments would obscure this continuity and potentially mislead interpretation, so we retain them to clearly illustrate the impact of SRC throughout the entire domain.
@@ IV Consequences in neutron stars with DM content
At this point, it is instructive to highlight the impact of incorporating DM through different formalisms. A direct comparison between Fig. 8 of the present work and Fig. 5 of Ref. [41] underscores the distinct modeling strategies adopted in the two studies. In Ref. [41], DM is treated within a single-fluid formalism and coupled to hadrons via Higgs boson exchange, whereas in the present work, we employ a two-fluid approach in which the dark and hadronic sectors interact exclusively through gravity. Although both studies include SRC and a quartic vector self-interaction term, the numerical difference in the constant $C$ is small and does not lead to significant modifications in the neutron star structure. Consequently, the main differences observed in the mass–radius relations arise from the distinct DM EoSs, the corresponding TOV formulations (single-fluid versus two-fluid), and the choice of RMF parametrization. In summary, these comparisons indicate that the effects of SRC remain robust across both frameworks, i.e., the increasing of the NS mass, while the detailed impact of DM is strongly dependent on the adopted EoS and fluid formalism.
@@
Fig. 10 illustrates how the mass of the last stable star, as determined from Figs. 8 and 9, varies as a function of the DM mass fraction $F_{\text{dm}}$.
@@
From Fig. 10, we compute $\Delta M$, defined as the difference between the maximum stellar mass obtained from the model with and without SRC. For the case $C\neq 0$, where $\Delta M>0$, we find $0.13<\Delta M/M_{\odot}<0.15$ across the range of $F_{\mathrm{dm}}$ considered. This interval is larger than that reported in Ref. [41]. In that study, a one-fluid formalism was employed to describe DM-admixed NSs, with the DM content controlled by the Fermi momentum of the dark particle ($k_{F}^{\mathrm{dm}}$). The hadronic sector was modeled by a relativistic framework including a quartic vector self-interaction term, and an increase in the maximum stellar mass was also observed when SRC effects were taken into account. From Fig. 8 of Ref. [41], we obtain the corresponding range in $\Delta M$ given by $0.059<\Delta M/M_{\odot}<0.066$ for $0<k_{F}^{\mathrm{dm}}<0.06~\mathrm{GeV}$.
@@ VI.3 General consequences
Figure 11: Neutron star mass-radius diagrams for two different parametrizations of the SRC model with $\eta_{p}\neq 0$ and $\eta_{n}\neq 0$ (full curves) and $\eta_{p}=\eta_{n}=0$. Crust EoS is not included.
At this point, we also emphasize the impact on the NS structure arising from the additional terms in the full expression of the chemical potentials when SRC are included. Since in the beta-equilibrated matter the condition $\mu_{n}-\mu_{p}=\mu_{e}=\mu_{\mu}$ must be satisfied, together with charge neutrality, $\rho_{p}=\rho_{e}+\rho_{\mu}$, the terms $2\rho_{n}\eta_{p}/(\pi^{2}\rho^{2})$ and $2\rho_{p}\eta_{n}/(\pi^{2}\rho^{2})$ modify the quantities determined from these equations, namely the proton fraction and the electron density. As a consequence, the total EoS used as input for solving the TOV equations (in either the one- or two-fluid formalism) differs from that obtained when the approximation is employed, as it depends on $y$ and $\rho_{e}$. In Fig. 11, we present the mass-radius relations for some parametrizations considered in this work, explicitly comparing the two cases.
@@ Acknowledgements.
This work has been done as a part of the Project INCT-Física Nuclear e Aplicações, under No. 408419/2024-5. It is also supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under Grants No. 444797/2024-6 (O.L., M.D.), No. 307255/2023-9 (O.L.), No. 301779/2025-2 (M.D.), No. 401565/2023-8 (Universal - O.L., M.D.), No. 409736/2025-2 (Universal - O.L., M.D.), and Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) under Thematic Project No. 2024/17816-8 (O.L., M.D.), and Project No 2025/06216-2 (C. B.).